OpenStudy (anonymous):

Solve the following absolute value equation: -4 |x + 5| + 3 = 21

OpenStudy (saifoo.khan):

piece of Pastry.

OpenStudy (anonymous):

haha

OpenStudy (saifoo.khan):

-4 |x + 5| = 18 |x + 5| = 18/-4 |x + 5| = -9/2 |x + 5| = -4.5 x = -4.5 - 5 & x = 4.5 -5 x = -9.5 & x = -0.5